40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14963     Entry time: Thu Oct 10 22:11:53 2019
Author: gautam 
Type: Update 
Category: LSC 
Subject: Trans QPD checkout 
  1. I removed the flip-mount that was installed on the EY in-air table for the mode-spectroscopy project (see Attachment #1). The Transmon QPD at EY sees IR light again.
  2. Dark noise checkout - see Attachment #2.
  3. Light-level expectations:
    • For the current config, let's say 0.8 W reaches the PRM, and we will have a PRG of 50. 
    • This implies ~5.5 kW circulating power in the arms.
    • This implies ~70mW will get transmitted through the ETM, of which at most half makes it to the QPD. 
    • In the nominal operating condition, we expect more like 6 W circulating in the arm cavity. So something like 30uW is expected to make it out onto the Trans QPDs.
    • But in this condition, we expect to run with the high-gain Thorlabs PD.
    • In reality the number is likely to be somewhat smaller. But we should set the transimpedance gain of this photodiode accordingly. Currently, there are a bunch of ND filters installed on this photodiode, which probably should be removed.
  4. Angular control
    • The other purpose these QPDs are expected to serve is to stabilize the angular motion of the cavities when locked with high circulating power.
    • Need to calculate what the sensing noise requirement is.
Attachment 1: EY_table_20191010.jpeg  6.321 MB  Uploaded Thu Oct 10 23:12:57 2019  | Hide | Hide all
EY_table_20191010.jpeg
Attachment 2: darkNoise.pdf  146 kB  Uploaded Thu Oct 10 23:14:19 2019  | Hide | Hide all
darkNoise.pdf
ELOG V3.1.3-